# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

### v2

@@ 6.3.1. Is Find-$\epsilon$-from-PRI practical?
  • Q5: SELECT SUM(fnlwgt) FROM adult WHERE capital_gain>0 AND income==’<=50K’ AND occupation==’Sales’
@@
Parameters. We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,\\ 0.099,\cdots,0.001]$ (37 values in total). We set $p=50$. We used six parallel processes for computing the PRIs.
@@
Figure 9. Average runtime (in log scale) to find $\epsilon$ over 10 runs
@@
Results. We first note that all queries on all 10 runs were able to find a suitable $\epsilon$ on every dataset we tested. Figure 9 shows a line plot where the x-axis is the data size and the y-axis is the runtime, both in log scale. Each line represents a query, with the error band showing the standard deviation.
As shown in the figure, the runtime increases linearly with the data size. Most queries took less than a few minutes to finish, even for 1 million records, except Q5, which took more than an hour. The runtime for Q5 is slower since it involves the attribute fnlwgt, which is a numerical attribute with a large domain size and many unique values, thus the query involves significantly more unique records than the other queries. For example, for 1 million records, Q5 has approximately 100000 unique records of which the PRIs need to be computed; whereas for other queries, the algorithm only needs to compute PRIs for no more than a few thousand records and the rest can be directly retrieved using memoization.
6.3.2. Effect of parameter $p$
The Find-$\epsilon$-from-PRI algorithm finds a suitable $\epsilon$ by letting controllers specify their privacy preferences through a parameter $p$, and checks whether $PRI_{min}/PRI_{max}\geq 1-(p/100)$ for each $\epsilon$. Smaller $p$ means stricter privacy preferences because the minimum and maximum PRI need to be closer. We test the effect of varying $p$ in the algorithm; specifically, when $p=5,25,50,75,95$.
Setup. We use the Adult dataset with 100000 rows that we generated in Section 6.3.1. We ran each of the five queries from Section 6.3.1 for 50 times, and report the $\epsilon$ chosen by our algorithm in each run.
@@ 6.3.2. Effect of parameter $p$
Figure 10. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$
@@
Results. Figure 10 shows boxplots of $\epsilon$ found by the algorithm for different $p$. As shown in the figure, setting larger $p$ tends to find larger $\epsilon$. Setting $p=5$ results in very small $\epsilon$ chosen for all five queries because the maximum PRI can only be no more than 5% from the minimum PRI; most $\epsilon$ values are less than 0.1 for Q1–4 and less than 1 for Q5. Queries with larger sensitivity (Q2, Q4, Q5) can find larger $\epsilon$ when $p>50$, and queries with sensitivity of 1 (Q1, Q3) tend to find smaller $\epsilon$. Overall, the results empirically demonstrate that the algorithm is useful for finding a suitable $\epsilon$ (and corresponding query output) that is consistent with the controller’s privacy preferences.
@@ 7. Related Work
Defining PRI. The PRI shares similar intuition as Personalized Differential Privacy (PDP) (Jorgensen et al., 2015) in that different individuals may need different levels of protection since some records are more sensitive than others. In contrast to PDP, which requires each individual to specify their own privacy requirement through a privacy budget, the PRI is computed by controllers to indicate the realized privacy loss of each individual if a DP output were to be released.
@@
Choosing privacy loss budget. Different approaches (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Hay et al., 2016; Ligett et al., 2017) aim to help users choose $\epsilon$ when using DP. One line of work (Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021) implements interfaces to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$. Nanayakkara et al. visualizes the disclosure risk in relation to $\epsilon$ using the definition of disclosure risk from Lee and Clifton, which is the probability of disclosing the presence or absence of any individual in the database. John et al. defines and visualizes the data sharing risk; Both metrics quantify potential data leakage, but neither quantifies leakage with respect to each individual in the data, unlike our PRI. Ligett et al. empirically finds the smallest $\epsilon$ consistent with an accuracy requirement in the private empirical risk minimization setting. Ge et al. translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound. Hsu et al. proposes an economic model for choosing $\epsilon$ by balancing the interests of the data analyst and the prospective participant who will participate if the benefit outweighs the risk. However, this model assumes that the participant can quantitatively express the ”cost” of whether to participate in the study, which is impractical for non-experts. Cummings et al. and Xiong et al. investigate users’ understanding of DP guarantees through multiple user studies, which offer insights on how to communicate the privacy implications of DP to users. To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
Interpreting utility. Tao et al. proposed DPXPlain to explain aggregate query answers by comparing answers with and without DP and assess whether the difference is due to the noise introduced by DP or properties of the data, thus helping analysts interpret whether the answers are useful for their analysis. This idea, complementary to ours, relies on analysts to come up with comparison questions that require explanation in order to understand any anomalies or trends in the query answers.
CDP LDP Shuffler
Crypto
Escrow
T
No Yes Yes Yes
Yes
Add.
Trust
Asm.
Trust
individuals
add noise
correctly
Trust
shuffler
Trust
crypto
impl.
Trust
SEV &
escrow
impl.
Encr.
Data?
No
Does not
collect
raw data
Does not
collect
raw data
Yes Yes
Error $O(1)$ $O(\sqrt{n}$) Near $O(1)$ $O(1)$ $O(1)$
Table 1. Comparisons of different DP deployment scenarios. T stands for trustworthy.
Deploying differential privacy. The threat model of the central DP model assumes adversaries cannot breach the platform and force unintended data flows, which is unrealistic. Alternatives beyond the central model aim to deploy DP with different trust assumptions. We summarize those different deployment scenarios in Table 1. The local model (LDP) has been proposed to eliminate the need for a trusted platform and it has been adopted in various real-world applications for collecting client data from their software and devices  (Erlingsson et al., 2014; Ding et al., 2017; Differential Privacy Team, 2017). In LDP, each individual perturbs their own data using LDP mechanisms before sending it to the platform; adversaries can control the platform but they only gain access to DP data. However, LDP also introduces a much larger error ($O(\sqrt{n})$, where n is the number of individuals) than the central DP model ($O(1)$). The shuffle model creates an anonymous communication channel that amplifies the privacy guarantees  (Kasiviswanathan et al., 2011; Balle et al., 2018), thus it requires adding less noise when users perturb their data and it offers better accuracy than the local DP model. However, it requires additional trust assumptions as the shuffler must be trusted. Lastly, techniques that leverage cryptographic primitives, specifically secure multiparty computation, homomorphic encryption, oblivious RAM, and trusted hardware enclaves, have been proposed to offer security guarantees when deploying DP without a trusted data curator  (Dwork et al., 2006a; Narayan and Haeberlen, 2012; Bater et al., 2016; Agarwal et al., 2018; Roth et al., 2019; Roy Chowdhury et al., 2020; Wang et al., 2022; Bater et al., 2018; Nguyen et al., 2020; Bittau et al., 2017; Bogatov et al., 2021). We showed that the escrow-based solution we introduce presents an interesting alternative with a lower computational cost.
@@ 8. Conclusions
In this paper, we showed that by including individual-level information about privacy loss (the PRI), controllers gain additional understanding of the effects of applying DP to the results of an analyst’s query. This translates into less burden choosing $\epsilon$: we implement this idea via an algorithm, Find-$\epsilon$-from-PRI.
In addition, we show that it is possible to augment analysts’ understanding of the effects of DP on their results by using a utility signaling protocol that reveals the isolated impact of DP on a downstream task. We then gather the requirements gleaned from the previous two contributions and propose a platform, based on a data escrow, to deploy both ideas.

### v3

@@ 6.2. RQ2: Are Both Algorithms Practical?
  • Q5: SELECT SUM(capital_gain) FROM adult
@@
Parameters. We initialize $\epsilon$ to test from 11 to 0.001 in fixed intervals; specifically, we test $\epsilon$ in $[10,\cdots,1,0.9,\cdots,0.1,0.09,\cdots,0.01,\\ 0.099,\cdots,0.001]$ (37 values in total). We set $p=5$ for Find-$\epsilon$-from-
PRI
, and $\epsilon_{svt}=1,\tau=10^{-6}$ for Find-and-release-$\epsilon$-from-PRI. We used eight parallel processes for computing the PRIs.
Configurations. We ran the experiments on a Chameleon Cloud (cha, [n. d.]) instance with 160 CPUs (Intel Xeon Platinum 8380) and 256GB memory. We used Ubuntu 22.04 and Python 3.10 for all experiments.
@@
Figure 8. Average runtime (in log scale) to find $\epsilon$ over 10 runs by Find-$\epsilon$-from-PRI (a) and Find-and-release-$\epsilon$-from-PRI (b)
@@
Results. Figure 8 shows two line plots where the x-axis is the data size and the y-axis is the runtime, both in log scale; the left plot corresponds to the run time of Find-$\epsilon$-from-PRI, and the right plot corresponds to the run time of Find-and-release-$\epsilon$-from-PRI. Each line in the plot represents a query, with the error band showing the standard deviation.
As shown in both plots, the runtime of both algorithms increases linearly with the data size. Most queries took less than a few minutes to finish, even for 1 million records. The runtime of Find-and-release-$\epsilon$-from-PRI is slightly longer than that of Find-$\epsilon$-from-PRI, especially when the dataset size is small. This is because applying SVT in Find-$\epsilon$-from-PRI takes an extra overhead since it needs to normalize the PRIs, compute the variance, and generate two Laplacian noise. However, when the dataset gets larger, the overhead becomes negligible and the runtime is dominated by the time to compute PRIs.
6.3. Microbenchmarks
@@
Figure 9. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$. The mechanism used is Laplace.
Figure 10. $\epsilon$ found by Find-$\epsilon$-from-PRI algorithm over 50 runs, with $p=5,25,50,75,95$. The mechanism used is Gaussian.
6.3.1. Effect of parameter $p$ in Find-$\epsilon$-from-PRI
The Find-$\epsilon$-from
-PRI
algorithm finds a suitable $\epsilon$ by letting controllers specify their privacy preferences through a parameter $p$, and checks whether $PRI_{min}/PRI_{max}\geq 1-(p/100)$ for each $\epsilon$. Smaller $p$ means stricter privacy preferences because the minimum and maximum PRI need to be closer. We test the effect of varying $p$ in the algorithm; specifically, when $p=5,25,50,75,95$.
Setup. We use the Adult dataset with 100000 rows that we generated in Section 6.2. We ran each of the five queries from Section 6.2 50 times, and recorded the $\epsilon$ chosen by the algorithm in each run.
@@ 6.3.1. Effect of parameter $p$ in Find-$\epsilon$-from-PRI
Results. Figure 9 and Figure 10 show the boxplots of $\epsilon$ found by the algorithm with different $p$ for each query; Figure 9 shows the results when the DP mechanism $\mathcal{M}$ used in the algorithm (for computing the query output $r$ and the PRIs) is Laplace Mechanism, and Figure 10 shows the results for the Gaussian mechanism. As shown in both figures, setting larger $p$ tends to find larger $\epsilon$. Setting $p=5$ results in very small $\epsilon$ chosen for all five queries because the maximum PRI can only be no more than 5% from the minimum PRI; most $\epsilon$ values are less than 0.1 when the mechanism is Laplace, and less than 1 when the mechanism is Gaussian. In fact, for most of the queries and $p$, using Gaussian caused the algorithm to choose larger $\epsilon$ compared to using Laplace, because the Gaussian Mechanism tends to produce noisier output than the Laplace Mechanism  [^5] This holds when the query’s $\ell_{2}$ sensitivity (used by Gaussian Mechanism) is similar to its $\ell_{1}$ sensitivity (used by Laplace Mechanism); otherwise, if the query’s $\ell_{2}$ sensitivity is much smaller, Gaussian Mechanism adds less noise compared to Laplace Mechanism. using the same $\epsilon$. In addition, both algorithms find larger $\epsilon$ when $p>50$ for queries with larger sensitivity (Q2, Q4, Q5), and they tend to find smaller $\epsilon$ for queries with sensitivity of 1 (Q1, Q3).
Overall, the results empirically demonstrate that both algorithms are useful for finding a suitable $\epsilon$ that is consistent with the controller’s privacy preferences.
Figure 11. $\epsilon$ found by Find-and-release-$\epsilon$-from-PRI algorithm over 50 runs, with $\tau=10^{-3},10^{-6},10^{-9},10^{-12}$.
6.3.2. Effect of parameter $\tau$ in Find-and-release-$\epsilon$-from-PRI
The Find-and-release-$\epsilon$-from-PRI algorithm lets controllers specify the PRI variance threshold $\tau$ (and the SVT budget $\epsilon_{svt}$), and it uses SVT to choose the first $\epsilon$ of which the variance of its corresponding PRIs is approximately smaller than the threshold. Smaller $\tau$ means stricter privacy preferences. We test the effect of varying $\tau$ in the algorithm; specifically, when $\tau=10^{-3},10^{-6},10^{-9},10^{-12}$.
Setup. We use the same setup as in Section 6.3.1. We set $\epsilon_{svt}=1$ and we use Laplace Mechanism as the DP mechanism $\mathcal{M}$.
Results. Figure 11 shows the boxplots of $\epsilon$ found by the algorithm with different $\tau$ for each query. The $\epsilon$ values selected are virtually similar and small when $\tau=10^{-6},10^{-9},10^{-12}$. On the contrary, the algorithm mostly selects the largest $\epsilon$ (10) it is considering when $\tau=10^{-3}$ for queries Q1, Q2, Q4, and Q5, and it selects a relatively small $\epsilon$ that is less 1 for Q3. This suggests controllers may need to observe the PRIs to decide a reasonable variance threshold to release $\epsilon$ and the query output $r$. As we noted before in Section 5.1, controllers may also rerun the algorithm multiple times with different parameters $\tau$ and $\epsilon_{svt}$ to observe how much noise is added and decide what to release that conforms to their privacy preferences.
@@ 7. Related Work
Privacy loss accounting. Various DP definitions have been proposed to account the privacy loss in a data-dependent manner or in different granularity. Our paper defines PRI as a relative indicator of the ex-post per-instance privacy loss (Redberg and Wang, 2021), which is built upon previous works that separately proposed ex-post privacy loss (Ligett et al., 2017) and per-instance privacy loss (Wang, 2019); ex-post privacy loss is the privacy loss conditioned on the realized output of a DP mechanism, and per-instance privacy loss is the privacy loss conditioned on a fixed dataset and a specific record in it. In addition,  Feldman and Zrnic (Feldman and Zrnic, 2021) proposed a privacy filter under Renyi differential privacy, which tracks the personalized privacy loss for each individual and drops an individual’s record when their personalized privacy loss exceeds a fixed budget (for all individuals). Seeman et al. (Seeman et al., 2024) proposed a formulation of DP called Per Record Differential Privacy; in this formulation, only the policy function on each individual’s privacy loss is published but not the actual privacy loss value, and individuals can compute their own privacy loss using the function. Unlike most works which define and use a different DP formation than standard DP, our work does not try to replace standard DP — we simply use the notion of ex-post per-instance privacy loss as a tool to connect our PRI definition to privacy loss in a more fine-grained granularity, and both our algorithms, Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI, satisfy $(\epsilon,\delta)$-DP.
@@
Choosing privacy loss budget. Different approaches (Hsu et al., 2014; Ge et al., 2019; Dwork et al., 2019; Lee and Clifton, 2011; Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021; Hay et al., 2016; Ligett et al., 2017) aim to help users choose $\epsilon$ when using DP. One line of work (Murtagh et al., 2018; Nanayakkara et al., 2022; Hay et al., 2016; Thaker et al., 2020; John et al., 2021) implements interfaces to visualize the effect of different variables (e.g., utility metrics) when adjusting $\epsilon$. Nanayakkara et al. (Nanayakkara et al., 2022) visualizes the disclosure risk in relation to $\epsilon$ using the definition of disclosure risk from Lee and Clifton (Lee and Clifton, 2011), which is the probability of disclosing the presence or absence of any individual in the database. John et al. defines and visualizes the data sharing risk; Both metrics quantify potential data leakage, but neither quantifies leakage with respect to each individual in the data, unlike our PRI. Ligett et al. (Ligett et al., 2017) empirically finds the smallest $\epsilon$ consistent with an accuracy requirement in the private empirical risk minimization setting. Ge et al. (Ge et al., 2019) translates a user-specified accuracy bound to a suitable DP mechanism, along with an $\epsilon$ that incurs the least privacy loss while satisfying the accuracy bound. Hsu et al. (Hsu et al., 2014) proposes an economic model for choosing $\epsilon$ by balancing the interests of the data analyst and the prospective participant who will participate if the benefit outweighs the risk. However, this model assumes that the participant can quantitatively express the ”cost” of whether to participate in the study, which is impractical for non-experts. Cummings et al. (Cummings et al., 2021) and Xiong et al. (Xiong et al., 2020) investigate users’ understanding of DP guarantees through multiple user studies, which offer insights on how to communicate the privacy implications of DP to users. To the best of our knowledge, no known approach helps users understand the implications of $\epsilon$ on individuals’ privacy, and use that to design an algorithm that finds $\epsilon$ automatically.
Customizing Privacy. Some works study methods to allow users to specify privacy requirements at the individual level. Jorgensen et al. (Jorgensen et al., 2015) proposes Personalized Differential Privacy in which each individual specifies their own privacy preferences instead of using a global privacy loss budget. Since choosing a numerical privacy preference value is challenging for typical users, the authors recommended associating suitable privacy preference values to different groups (e.g., low/medium/high privacy). He et al. (He et al., 2014) proposes Blowfish Privacy in which in addition to specifying $\epsilon$, individuals specify their privacy policies consisting of secrets, i.e., information that must be kept secret about the individuals, and constraints, i.e., information that may be known publicly and exploited by adversaries. Both works attempt to define the granularity of privacy protection on the individual level but put the burden on the users to specify their privacy preferences. Instead, PRI is interpretable and controllers can specify their privacy preferences by observing the PRI of each individual directly.
@@ 8. Conclusions
In this paper, we showed that by including individual-level information about privacy loss (the PRI), controllers gain more understanding of the effects of applying differential privacy to the output of an analyst’s query. This translates into less burden in choosing $\epsilon$: we implement this idea via an algorithm, Find-$\epsilon$-from-PRI. In addition, we propose an alternative algorithm, called Find-and-release
-$\epsilon$-from-PRI
, which finds and releases $\epsilon$ privately. We present solutions to bound privacy loss when answering multiple queries without requiring controllers to set a global privacy budget. Overall, our work reduces the burden for controllers to use DP in practice.